Would you please help me find :
Which of the lines has the same slope and passes through point (6,12) ?
A. Y-2X=0
B. Y+3X=30
C. Y-3X= -6
D. Y+3X=42
I appreciate your help. Jose

Answer by Alan3354(69443)

Which of the lines has the same slope and passes through point (6,12) ?
A. Y-2X=0
B. Y+3X=30
C. Y-3X= -6
D. Y+3X=42
===============
B & D have the same slope.
Sub 6 for x and 12 for y to see if they pass thru the point.